OpenStudy (anonymous):

I have a rectangle that I've measured at 15" across at opposite corners. What is the width and height?

jimthompson5910 (jim_thompson5910):

You would need more info because there are infinitely many rectangles that have a diagonal of 15 inches.

jimthompson5910 (jim_thompson5910):

If you knew how the length and width are connected (eg: the length is 3 times the width), then you can use that to find the answer

OpenStudy (anonymous):

It's a computer monitor. I just have the diagonal of 15 inches :(

jimthompson5910 (jim_thompson5910):

Does it give you a ratio of the length to width, like 4:3 or 16:9 ?

OpenStudy (anonymous):

Best guess: 16:9

jimthompson5910 (jim_thompson5910):

ok so that means that the width is 16 inches and the height is 9 inches OR the width is 16k inches and the height is 9k inches for some positive number k

jimthompson5910 (jim_thompson5910):

So this means a^2 + b^2 = c^2 L^2 + W^2 = 15^2 (16k)^2 + (9k)^2 = 225 256k^2 + 81k^2 = 225 337k^2 = 225 k^2 = 225/337 k^2 = 0.667655786 k = sqrt(0.667655786) k = 0.8171020 Therefore, Width: 16k = 16(0.8171020) = 13.073632 inches (roughly) Height: 9k = 9(0.8171020) = 7.353918 inches (roughly)

jimthompson5910 (jim_thompson5910):

This is of course all hinged on the assumption that the monitor is 16:9

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

yw